assume that we want to seat Alex, Bonnie, Carl, Daria, Edith, and Frank in a row of six chairs. In how many ways can the seating be done if Alex and Bonnie must sit together.
Answer by Edwin McCravy(20054)

It's the number of permutation of the 5 things {AB,C,D,E,F}
plus the number of permutations of the 5 things {BA,C,D,E,F}

Answer 5!+5! = 2×5! = 2×120 = 240 ways.

Edwin